HUBBLE SPACE TELESCOPE DAILY REPORT #2209 PERIOD COVERED: 0000Z (UTC) 09/15/98 - 0000Z (UTC) 09/16/98 Daily Status Report as of 259/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed NIC/2 7887 (A Survey of Gravitational Lenses as Cosmological Tools II) The Near Infrared Camera (Camera Two) was used to make observations of the gravitational lens candidate system HST14164+5. The purpose is to extend our IR survey of gravitational lenses and gravitational lens candidates by obtaining NICMOS images of 14 additional lens systems or candidates. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Eight Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/2 7895 (HST NICMOS SNAPSHOT Observations of Field Galaxies at Z < 1) The Near Infrared Camera (Camera Two) was used to make observations of the field galaxy Gal-141746. The proposed NICMOS observation will provide their morphology in rest frame near infrared (NIR) wavelengths {8000 - 16000 A}, where the effect of dust extinction is less severe, and their NIR colors. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed STIS/CCD 7930 (STIS CCD Binned Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make numerous bias exposures in 1 x 2 binned mode in order to permit construction of "superbias" frames in which the effects of read noise have been rendered negligible. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.7 Completed Two Sets of NIC/1/2/3 7898 (Snapshot Survey of Proto-planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the protoplanetary nebula IRAS22223 and around star HD179821. This is to help discern the structure of the nebulae and the evolutionary sequence that produces axially symmetric planetary nebulae. As described in HSTAR 6674 and yesterday's report, during the second iteration of this proposal, the guide star acquisition failed and the take data flag remained down for one observation. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.8 Completed Two Sets of STIS/MA1 7439 (H-Lyman Alpha Emission from the Upper Atmosphere of Uranus) The Space Telescope Imaging Spectrograph (MA1) was used to make H-Lyman Alpha observations of the emissions from the upper atmosphere of the planet Uranus. An excess of H-Lyman Alpha emission from Uranus' sunlit hemisphere has been detected by the IUE satellite since 1982, and some excess was confirmed with the Voyager 2 UVS during the 1986 encounter with Uranus. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star LHS511. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/MA2 7720 (MAMA Full Field Sensitivity Monitor and PSF Check) The Space Telescope Imaging Spectrograph (MA2) was used to monitor the sensitivity of the FUV-MAMA and NUV-MAMA over the full field and monitor the PSF. This was done by making observations of NGC-6681. The observations completed as planned, with no reported problems. 1.12 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 6794 (Proper Motions of Herbig-Haro Jets) The WF/PC-2 was used to make observations of the Herbig-Haro object HH34-P1 in order to measure the proper-motions of these high Mach number flows of its two opposing jets. The tangential velocities of several hundred km/s will produce displacements between our first- and second-epoch images of 2-4 pixels after only two years. The observations were completed as planned, and no anomalies were reported. 1.14 Completed Four Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed Two Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.17 Completed Two Sets of WF/PC-2 6735 (Locating Disks, Accretion Flows, and outflows in Close Binary T Tauri Stars) The WF/PC-2 was used to make observations of the binary system, T-Tauri class star FY-Tau in order to study the location of disks and accretion flows. Although the standard model for T-Tauri star formation is based on single stars, we know that these T Tauri stars are in multiple star systems having separations less than 100 AU. The observations were completed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The SSR EDAC error counter was reset three times per ROP SR-1. The HST IP was the prime system again from approximately 258/1200Z until approximately 258/2000Z. There were no operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH